the equivalent expression of sin300° is - sin____°
Found 2 solutions by Boreal, Theo:
Answer by Boreal(15235)

-sin 60.
sin 270 is -1, and 300 degrees is the equivalent of 60 degrees here.

Answer by Theo(13342)

300 degrees is in the fourth quadrant because it is between 270 and 360 degrees.

the equivalent angle in the first quadrant is equal to 360 - 300 = 60 degrees.

the equivalent angle in the second quadrant is 180 - 60 = 120 degrees.

the equivalent angle in the third quadrant is 180 + 60 = 240 degrees.

the sine function is positive in the first and second quadrant.

the sine function is negative in the third and fourth quadrant.

since sine(300) is in the fourth quadrant, then its sine will be negative.

the equivalent angle will be the sine function in the third quadrant.

your equivalent angle is the refore 240 degrees.

sin(300) = -.8660254038
sin(240) = -.8660254038

they're the same.

in the first and second quadrant, the sine function will be the same value but will be positive.

sin(60) = .8660254038
sin(120) = .8660254038

the value of the sine function is the same in all 4 quadrant, except that:

in the first and second quadrants, it is positive.
in the third and fourth quadrants, it is negative.

when you use your calculator to get the sine of the angle, make sure your calculator is in degree mode.

otherwise you'll get some off the wall answers that won't agree with what is shown above.